OpenStudy (anonymous):

- 2/3 y = - 4/15 What are the steps to solve?

OpenStudy (lgbasallote):

you can start by multiplying both sides by 3

OpenStudy (anonymous):

I would multiply both sides by 15. This will eliminate the denominators and give you a simpler set of numbers to work with.

OpenStudy (anonymous):

Discard the negative sign on each side. Multiply each side by 15, you get: \[ 10y = 4 \rightarrow y = \dfrac {4}{10} = \boxed {\dfrac {2}{5}}. \]

OpenStudy (anonymous):

How do I know when to discard negative signs?

OpenStudy (anonymous):

When there is a negative on both sides, you can discard it - that's a colloquial way of saying "multiply each side by -1".

OpenStudy (anonymous):

oh ok so -2/3 y = -4/5 is equivalent to 2/3 y = 4/5 makes sense.

OpenStudy (anonymous):

Thanks for the help, much appreciated.